OpenStudy (anonymous):

use de moivre's formula with n= 2 to find trigonometric identities cos2theta and sin 2theta plz help me.....:(

OpenStudy (anonymous):

farily simple

OpenStudy (anonymous):

fairly lol

OpenStudy (anonymous):

\[(\cos(\theta) + isin(\theta))^n = cos(n \theta) +isin(n \theta ) ) \]

OpenStudy (watchmath):

\(\cos (2\theta)_1\sin(2\theta)=(\cos \theta +i \sin \theta)^2=(\cos^2\theta-sin^2\theta)+i(2\sin\theta\cos\theta)\)

OpenStudy (anonymous):

^ that yeh :|

OpenStudy (watchmath):

The left most expression should be \(\cos(2\theta)+i\sin(2\theta)\)

OpenStudy (anonymous):

there is actually something to do here yes? besides using laws of exponents i mean.

OpenStudy (anonymous):

oops my apologies whatchmath wrote it. should should also mumble something about setting real part = real part and imaginary part = imaginary part.

OpenStudy (anonymous):

watchmath i mean. yikes it is early

OpenStudy (watchmath):

Are you in US too satellite?

OpenStudy (anonymous):

yes US

OpenStudy (anonymous):

trying to learn latex and practicing here as much as i can